# Bipartite Graph Variational Auto-Encoder with Fair Latent Representation to Account for Sampling Bias in Ecological Networks

### v2

@@ 6 Conclusion
In this paper, we proposed not only a bipartite extension of the graph variationnal auto-encoder, but also a new method to have a fair latent representation with respect to a continuous variable. We handled a particular structure of data: observed collections of pollination events with numerous covariates from which are derived a plant-pollinator interaction network. We then used our model to tackle the sampling effects affecting this large ecological network, an issue of particular importance when the sampling involves citizen science. Even if our proposed model is specific to the study of Spipoll, our contribution using GNN and the HSIC as a bias mitigation strategy can be adapted to provide interesting results in other fields of study where both networks and sampling bias are at stake. A question of interest that we left for future research is to assess the effects of the covariates on the pollination interactions.
@@ Acknowledgments
Appendix A Proof of proposition 1
@@
$\frac{\partial L}{\partial\lambda_{1}}=1-||v||^{2}=0,\quad\frac{\partial L}{\partial\lambda_{2}}=||S^{\top}Xv||^{2}=0.$
@@ Appendix B Extensive simulation study
All the simulations are reproducible and available on
https://github.com/AnakokEmre/Bipartite-and-fair-VGAE.
@@ B.1.1 Setting
First, we fit a classical bipartite and variational graph auto-encoder on $B_{i,j}$. We expect that this auto-encoder would yield a latent representation $\tilde{Z_{1}}$ correlated with $S$ and $T$. We then fit our bipartite and fair auto-encoder to compare the results and see if the yielded latent space is independent of $S$. We also compare our methodology with an adversarial learning algorithm (ADV) (Zhang et al., 2018) where the output $\mu_{1}$ is then used as an input to a 4-layer perceptron, which attempts to predict the protected variable $S$. The loss is then penalized if the predicted output is correlated with the protected variable.
@@ B.1.2 Results
The results for the link prediction task in the simulated network are summarized in Table 3. The simulations were done with dataset splits, with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. We compare the methods with the area under the ROC curve (AUC) score, the $HSIC$ between the latent space $\tilde{Z_{1}}$ and $S$, the number of times the p-value associated with the HSIC independence test is lower than $0.05\%$ (#$\#p_{0.05}$) and the Euclidean norm of the correlation matrix between $\tilde{Z_{1}}$ and $S$ ($cor$). In the table, are reported the mean and standard deviation for 100 trials, except for #$\#p_{0.05}$ which is only a count. We set the hyperparameter $\delta=n_{1}$. For each trial, the simulations begin with 10 random initializations, and were fit using 1000 iterations of the Adam algorithm with learning rate $0.01$, using a computer equipped with an Intel Xeon(R) CPU E5-1650 v4 and 32GB of RAM. The model that achieved the most favorable performance on the validation test set is then selected to evaluate the performance on the test dataset.
@@
An example of the latent space of BVGAE and fair-BVGAE can be seen in figure 6. Looking at the coloring, we can see for the BVGAE that the latent space is clearly correlated with $S$ , while the latent space of the fair-BVGAE does not share structure with the protected variable $S$. The HSIC test between the fair latent space and $S$ yields us a p-value equals to $0.139$, we do not reject the hypothesis that the latent space $Z_{1}$ is independent of $S$. Simulation with binary protected variable is available in Section B.3.
@@ B.2 Impact of hyperparameter $\delta$
We remind the expression of the variational loss given in Section 3:
@@
In this expression, $\delta$ is the hyperparameter associated with the $RFF\>HSIC$. Setting $\delta=0$ yields the same result as fitting the classical BVGAE. The following simulation study is performed to study the impact of this hyperparameter on the different scores.
@@ B.2.1 Setting
The settings are nearly identical as in Section B.1. In this simulation, we are going to generate bipartite networks made of $n_{1}=1000$ rows and $n_{2}=100$ columns. Let $S_{i}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$ for $i=1,\dots,n_{1}$ and $T_{i}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$ for $i=1,\dots,n_{1}$ and independent of $S$. We suppose that $S$ is the protected variable. Let $Z_{1}=(S,T)\in\mathbb{R}^{n_{1}\times 2}$ be the 2-column matrix made with both $S$ and $T$. Let $Z_{2}\overset{i.i.d.}{\sim}\mathcal{N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right)\in\mathbb{R}^{n_{2}\times 2}$. We simulate our bipartite adjacency matrix with Bernoulli $B_{i,j}\overset{i.i.d.}{\sim}\mathcal{B}(sigmoid(z_{1i}^{\top}\mathbf{I}_{D_{+},D_{-}}z_{2j}))$.
@@ B.2.2 Results
The results for the link prediction task in the simulated network are summarized in Table 4. The simulations were done with dataset splits, with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. In the table are reported the mean and standard deviation for 100 trials, except for #$\#p_{0.05}$ which is only a count.
@@
Figure 7: Impact of the parameter $\delta$ on the $AUC$ (upper left), the norm of the correlation matrix (upper right), the log10 HSIC (bottom left), and the p-value of the independence test (bottom right).
@@
±$\pm 1.22\times 10^{-2}$$\pm 1.56\times 10^{-4}$$\pm 4.37\times 10^{-6}$$\pm 2.72\times 10^{-6}$$\pm 1.65\times 10^{-6}$$\pm 1.27\times 10^{-6}$$\pm 1.50\times 10^{-6}$
#$\#p_{0.05}$ 100/100 44/100 0/100 0/100 0/100 0 /100 0/100
$cor$ $0.735\pm 0.206$ $0.035\pm 0.055$ $0.011\pm 0.007$ $0.011\pm 0.008$ $0.011\pm 0.008$ $0.011\pm 0.008$ $0.017\pm 0.025$
@@ B.3.1 Setting
Simulations with a similar setting as in Section B.1 has been performed with a simulated latent space structured along a binary protected variable $S\in\{-1,1\}$.
@@ B.3.2 Results
Results for the link prediction task in the simulated network are summarized in Table 5. The simulations were done with dataset splits, with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. In the table are reported the mean and standard deviation for 100 trials, except for #$\#p_{0.05}$ which is only a count. We set the hyperparameter $\delta=n_{1}=1000$. For each trial, the simulations begin with 10 random initialization, and were fit using 1000 iterations of the Adam algorithm with learning rate $0.01$. The model that achieved the most favorable performance on the validation test set is then selected to evaluate the performance on the test dataset.
@@
#$\#p_{0.05}$ 100/100 0/100 54/100
$cor$ $0.964\pm 0.020$ $0.014\pm 0.029$ $0.246\pm 0.322$
B.4 Comparison of fair linear embedding
We wish to perform a linear embedding
@@ B.4 Comparison of fair linear embedding
B.4.1 Principal component analysis using $X$
$\mathcal{L}(W_{0},W_{1})=\frac{1}{n}||g_{W_{1}}(f_{W_{0}}(X))-X||.$
B.4.2 Principal component analysis using $P_{S^{\bot}}X$
$\mathcal{L}(W_{0},W_{1})=\frac{1}{n}||g_{W_{1}}(f_{W_{0}}(P_{S^{\bot}}X))-X||.$
B.4.3 Principal component analysis using $X$ and the HSIC loss
$\mathcal{L}(W_{0},W_{1})=\frac{1}{n}||g_{W_{1}}(f_{W_{0}}(X))-X||+\delta RFF\>HSIC(f_{W_{0}}(X),S).$
Here we have chosen $\delta=10^{5}$.
Table 6: Comparison between the PCA, the projected PCA and the HSIC PCA
PCA PCA with projection PCA with HSIC
MSE $7.16\times 10^{-2}\pm 21.4\times 10^{-2}$ $9.88\pm 6.34$ $8.80\pm 5.54$
HSIC $3.15\times 10^{-2}$ $1.15\times 10^{-3}$ $2.21\times 10^{-3}$
±$\pm 1.02\times 10^{-2}$$\pm 0.60\times 10^{-3}$$\pm 4.81\times 10^{-3}$
#$\#p_{0.05}$ 100/100 3/100 10/100
$cor$ $0.991\pm 0.229$ $6.83\times 10^{-2}\pm 4.92\times 10^{-2}$ $0.142\pm 0.241$
B.5 Temporal gain using RFF HSIC
The purpose of this study is to see the temporal gain of using the RFF HSIC instead of $\widehat{HSIC}$ for different value of $n$, and to see if $RFF\ HSIC$ is an accurate approximation of $\widehat{HSIC}$.
B.5.1 Settings
For various value of $n$, we consider $X$ a $n\times 4$ matrix, with entries such as $X_{i,j}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Let $S$ be a $n\times 4$ matrix. Under the null hypothesis, we consider that $S$ is independent of $X$, with $S_{i,j}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Under the alternative hypothesis we consider that $S=3X$. The aim is to compute the HSIC between $X$ and $S$ with $\widehat{HSIC}(X,S)$, $RFF\ HSIC(X,S)$ and their respective gradient, under both hypothesis. In our fairness setting, $S$ would represent the protected variable and would be fixed once and for all, while $X$ would change according to the computed gradient. Therefore, we evaluate $L_{i,j}=L(s_{i},s_{j})=e^{-\frac{||s_{i}-s_{j}||^{2}}{2}}$, $L^{\prime}=\sum_{1\leq p,q\leq n}L_{q,p}$, and $L^{\prime\prime}_{i}=\sum_{q=1}^{n}L_{i,q}$ in advance to perform a quicker computation :
@@ B.5.1 Settings
For multiple random realisations of $X$, we compute $K_{i,j}=e^{-\frac{||x_{i}-x_{j}||^{2}}{2}}$ and then plug it in Equation 8. We also compute the gradient with respect to $X$ using Pytorch automatic differentiation, before measuring the average time taken to realize both of these operations.
Using the method presented in Equation 3, we also compute the $RFF\ HSIC$ between $X$ and $S$ with $D=\left\lceil\sqrt{n}\right\rceil$, and its gradient using Pytorch automatic differentiation. We measure the average time taken to calculate the $RFF\ HSIC$ and its gradient compared to the $\widehat{HSIC}$. We also compare how close the value of $RFF\ HSIC$ is to the $\widehat{HSIC}$ using the squared error between the two with $\widehat{HSIC}$ as a base value. All theses computation are realized on an Intel Xeon(R) CPU E5-1650 v4 and 32GB of RAM.
B.5.2 Results
Figure 11: Estimated time to compute $HSIC(X,S)$ (left) and its gradient with respect to $X$(right) using $\widehat{HSIC}$ and $RFF\ HSIC$.
Figure 12: Estimated squared error between the $RFF\ HSIC$ and the $\widehat{HSIC}$
As we can see in Figure 11, using the $RFF\ HSIC$ is much faster than $\widehat{HSIC}$ by a large margin. Under the null hypothesis, the estimation is less accurate than under the alternative hypothesis (Figure 12) but the hypothesis doesn’t affect the computation time. For 1000 iterations and for $n=10000$, the $RFF\ HSIC$ and its gradient would require around 7.9 seconds of computation time, while the $\widehat{HSIC}$ would require around 35 minutes. In the Spipoll dataset, we have considered $n=12574$ with a latent space of dimension 4. Using a second order polynomial, we can estimate that computing $1000$ times the HSIC and its gradient would require around 56 minutes for $\widehat{HSIC}$ and 11 seconds for $RFF\ HSIC$. We also only presented results from data in the time period between 2017 and 2020, but if we considered the Spipoll data set from 2010 to 2020, then $n\approx 26000$. In this case, we can estimate that computing $1000$ times the HSIC and its gradient would require around 4 hours with $\widehat{HSIC}$ while the $RFF\ HSIC$ would only need 33 seconds. All these estimations are done without taking into account the fact that the computation of the $n\times n$ Gram matrix, needed for the $\widehat{HSIC}$ can also require a lot of memory from the computer.
B.6 Spipoll, exploration with higher dimensional latent spaces
Figure 13: Estimated mean on 10 trials for the $AUC(\widehat{B})$ (left) and $AUC(\widehat{B^{\prime}})$ (right) for link prediction in the Spipoll data set using BGVAE (top) and the fair-BGVAE (bottom) for various values of $D_{+}$ and $D_{-}$.
In Section 5 we show in detail the results for the case where the latent space has 4 dimensions with $D_{+}=D_{-}=2$. We justify this choice by looking at the estimated mean of the $AUC(\widehat{B})$ for different number of dimension for the latent space. Looking at Figure 13, we can see that the $AUC(\widehat{B})$ doesn’t change for higher values of $D_{+}$ and $D_{-}$.
@@ Appendix C Latent space representation of Spipoll
C.1 Latent space representation
Figure 14: Observed initial plant-insect network
In addition to the plant-pollinator network, the method provides a session-pollinator embedding that can be seen in Figure 15 and Figure 16. We remind that $D_{+}=D_{-}=2$, which means that for the first two dimensions, insects and sessions that are embedded in the same direction are more likely to be connected, and the ones in the opposite direction are less likely to be connected. On the contrary, insects and sessions that are embedded in the same direction for the third and fourth dimensions are less likely to be connected, while the ones in the opposite direction are more likely to be connected. A closer look of the latent space highlights the ecological interest of such representation in Section C.2.

### v3

@@ 6 Conclusion
In this paper, we proposed not only a bipartite extension of the graph variational auto-encoder, but also a new method to have a fair latent representation with respect to a continuous variable. We handled a particular structure of data: observed collections of pollination events with numerous covariates from which are derived a plant-pollinator interaction network. We then used our model to tackle the sampling effects affecting this large ecological network, an issue of particular importance when the sampling involves citizen science. Even if our proposed model is specific to the study of Spipoll, our contribution using GNN and the HSIC as a bias mitigation strategy can be adapted to provide interesting results in other fields of study where both networks and sampling bias are at stake. A question of interest that we left for future research is to assess the effects of the covariates on the pollination interactions.
@@ Acknowledgments
Appendix A The linear embedding case
A.1
Let $X$ a $n\times d$ matrix and let $S$ be $n\times d_{s}$ matrix. Without loss of generality, we assume that each column of $X$ has been centered. We wish to perform a one dimensional principal component analysis on $X$ that would yield us a vector $v$ and a lower dimensional embedding of $X$ given by $Xv$ that maximizes the variance. However, we wish to have a latent representation $Xv$ independent of the protected variable $S$. If we were in the context of probabilistic PCA (tippingProbabilisticPrincipalComponent1999) where $X$ and $S$ would have been multivariate Gaussian, projecting $X$ onto the space orthogonal to $S$: $S^{\perp}$ beforehand would have been enough to guarantee the independence between $S$ and the latent representation $Xv$, this can be solved using PPCA with covariates (kalaitzisResidualComponentAnalysis2012). We show that this approach is equivalent to find the optimal projection with respect to an independence constraint.
@@ Appendix A The linear embedding case
[…]
@@
A.2 Comparison of fair linear embedding
We wish to perform a linear embedding
$X\xrightarrow[Linear]{}Z\xrightarrow[Linear]{}\widehat{X}$
A.2.1 Principal component analysis using $X$
[…]
A.2.2 Principal component analysis using $P_{S^{\bot}}X$
[…]
A.2.3 Principal component analysis using $X$ and the HSIC loss
[…]
Here we have chosen $\delta=10^{5}$.
Table 3: Comparison between the PCA, the projected PCA and the HSIC PCA
PCA PCA with projection PCA with HSIC
MSE $7.16\times 10^{-2}\pm 21.4\times 10^{-2}$ $9.88\pm 6.34$ $8.80\pm 5.54$
HSIC $3.15\times 10^{-2}$ $1.15\times 10^{-3}$ $2.21\times 10^{-3}$
$\pm 1.02\times 10^{-2}$ $\pm 0.60\times 10^{-3}$ $\pm 4.81\times 10^{-3}$
[…]
corcor $0.991\pm 0.229$ $6.83\times 10^{-2}\pm 4.92\times 10^{-2}$ $0.142\pm 0.241$
@@ Appendix B Extensive simulation study
All the simulations are reproducible and available on https://github.com/AnakokEmre/Bipartite-and-fair-VGAE.
@@ B.1.1 Setting
First, we fit a classical bipartite and variational graph auto-encoder on $B_{i,j}$. We expect that this auto-encoder would yield a latent representation $\tilde{Z_{1}}$ correlated with $S$ and $T$. We then fit our bipartite and fair auto-encoder to compare the results and see if the yielded latent space is independent of $S$. We also compare our methodology with an adversarial learning algorithm (ADV) (zhangMitigatingUnwantedBiases2018) where the output $\mu_{1}$ is then used as an input to a 4-layer perceptron, which attempts to predict the protected variable $S$. The loss is then penalized if the predicted output is correlated with the protected variable.
@@ B.1.2 Results
The results for the link prediction task in the simulated network are summarized in Table 4. The simulations were done with dataset splits, with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. We compare the methods with the area under the ROC curve (AUC) score, the $HSIC$ between the latent space $\tilde{Z_{1}}$ and $S$, the number of times the p-value associated with the HSIC independence test is lower than $0.05\%$ ($p_{0.05}$) and the Euclidean norm of the correlation matrix between $\tilde{Z_{1}}$ and $S$ (corcor). In the table, are reported the mean and standard deviation for 100 trials, except for $p_{0.05}$ which is only a count. We set the hyperparameter $\delta=n_{1}$. For each trial, the simulations begin with 10 random initializations, and were fit using 1000 iterations of the Adam algorithm with learning rate $0.01$, using a computer equipped with an Intel Xeon(R) CPU E5-1650 v4 and 32GB of RAM. The model that achieved the most favorable performance on the validation test set is then selected to evaluate the performance on the test dataset.
@@
An example of the latent space of BVGAE and fair-BVGAE can be seen in figure 8. Looking at the coloring, we can see for the BVGAE that the latent space is clearly correlated with $S$ , while the latent space of the fair-BVGAE does not share structure with the protected variable $S$. The HSIC test between the fair latent space and $S$ yields us a p-value equals to $0.139$, we do not reject the hypothesis that the latent space $Z_{1}$ is independent of $S$. Simulation with binary protected variable is available in B.3.
@@ B.2 Impact of hyperparameter $\delta$
We remind the expression of the variational loss given in Section 2.4:
@@
In this expression, $\delta$ is the hyperparameter associated with the RFFHSICRFF\>HSIC. Setting $\delta=0$ yields the same result as fitting the classical BVGAE. The following simulation study is performed to study the impact of this hyperparameter on the different scores.
@@ B.2.1 Setting
The settings are nearly identical as in B.1. In this simulation, we are going to generate bipartite networks made of $n_{1}=1000$ rows and $n_{2}=100$ columns. Let $S_{i}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$ for $i=1,\dots,n_{1}$ and $T_{i}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$ for $i=1,\dots,n_{1}$ and independent of $S$. We suppose that $S$ is the protected variable. Let $Z_{1}=(S,T)\in\mathbb{R}^{n_{1}\times 2}$ be the 2-column matrix made with both $S$ and $T$. Let $Z_{2}\overset{i.i.d.}{\sim}\mathcal{N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right)\in\mathbb{R}^{n_{2}\times 2}$. We simulate our bipartite adjacency matrix with Bernoulli $B_{i,j}\overset{i.i.d.}{\sim}\mathcal{B}(sigmoid(z_{1i}^{\top}\mathbf{I}_{D_{+},D_{-}}z_{2j}))$.
@@ B.2.2 Results
The results for the link prediction task in the simulated network are summarized in Table 5. The simulations were done with dataset splits, with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. In the table are reported the mean and standard deviation for 100 trials, except for $p_{0.05}$ which is only a count.
@@
Figure 9: Impact of the parameter $\delta$ on the AUCAUC (upper left), the norm of the correlation matrix (upper right), the log10 HSIC (bottom left), and the p-value of the independence test (bottom right).
@@
$\pm 1.22\times 10^{-2}$ $\pm 1.56\times 10^{-4}$ $\pm 4.37\times 10^{-6}$ $\pm 2.72\times 10^{-6}$ $\pm 1.65\times 10^{-6}$ $\pm 1.27\times 10^{-6}$ $\pm 1.50\times 10^{-6}$
[…]
corcor $735\pm 0.206 0.0$$35\pm 0.055 0.0$$1\pm 0.007 0.011$$\pm 0.008 0.011$$\pm 0.008 0.011$$1\pm 0.008 0.0$$017\pm 0.025$
@@ B.3.1 Setting
Simulations with a similar setting as in B.1 has been performed with a simulated latent space structured along a binary protected variable $S\in\{-1,1\}$.
@@ B.3.2 Results
Results for the link prediction task in the simulated network are summarized in Table 6. The simulations were done with dataset splits, with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. In the table are reported the mean and standard deviation for 100 trials, except for $p_{0.05}$ which is only a count. We set the hyperparameter $\delta=n_{1}=1000$. For each trial, the simulations begin with 10 random initialization, and were fit using 1000 iterations of the Adam algorithm with learning rate $0.01$. The model that achieved the most favorable performance on the validation test set is then selected to evaluate the performance on the test dataset.
@@
$p_{0.05}$ 100/100 0/100 54/100
corcor $964\pm 0.020 0.0$$014\pm 0.029$ $0.246\pm 0.322$
B.4 Temporal gain using RFF HSIC
The purpose of this study is to see the temporal gain of using the RFF HSIC instead of $\widehat{HSIC}$ for different value of $n$, and to see if RFFHSICRFF\ HSIC is an accurate approximation of $\widehat{HSIC}$.
B.4.1 Settings
@@
[…]
Using the method presented in Equation 2, we also compute the RFFHSICRFF\ HSIC between $X$ and $S$ with $h=\left\lceil\sqrt{n}\right\rceil$, and its gradient using Pytorch automatic differentiation. We measure the average time taken to calculate the RFFHSICRFF\ HSIC and its gradient compared to the $\widehat{HSIC}$. We also compare how close the value of RFFHSICRFF\ HSIC is to the $\widehat{HSIC}$ using the squared error between the two with $\widehat{HSIC}$ as a base value. All theses computation are realized on an Intel Xeon(R) CPU E5-1650 v4 and 32GB of RAM.
B.4.2 Results
Figure 12: Estimated time to compute $HSIC(X,S)$ (left) and its gradient with respect to XX(right) using $\widehat{HSIC}$ and RFFHSICRFF\ HSIC.
Figure 13: Estimated squared error between the RFFHSICRFF\ HSIC and the $\widehat{HSIC}$
As we can see in Figure 12, using the RFFHSICRFF\ HSIC is much faster than $\widehat{HSIC}$ by a large margin. Under the null hypothesis, the estimation is less accurate than under the alternative hypothesis (Figure 13) but the hypothesis doesn’t affect the computation time. For 1000 iterations and for $n=10000$, the RFFHSICRFF\ HSIC and its gradient would require around 7.9 seconds of computation time, while the $\widehat{HSIC}$ would require around 35 minutes. In the Spipoll dataset, we have considered $n=12574$ with a latent space of dimension 4. Using a second order polynomial, we can estimate that computing $1000$ times the HSIC and its gradient would require around 56 minutes for $\widehat{HSIC}$ and 11 seconds for RFFHSICRFF\ HSIC. We also only presented results from data in the time period between 2017 and 2020, but if we considered the Spipoll data set from 2010 to 2020, then $n\approx 26000$. In this case, we can estimate that computing $1000$ times the HSIC and its gradient would require around 4 hours with $\widehat{HSIC}$ while the RFFHSICRFF\ HSIC would only need 33 seconds. All these estimations are done without taking into account the fact that the computation of the $n\times n$ Gram matrix, needed for the $\widehat{HSIC}$ can also require a lot of memory from the computer.
@@ Appendix C Latent space representation of Spipoll
Figure 14: Observed initial plant-pollinator network
The observed plant-pollinator network is provided in Figure 14. In addition to the reconstructed plant-pollinator network, the method provides an embedding of dimension $D=D_{+}+D_{-}=4$ with $D_{+}=D_{-}=2$, which means that for the first two dimensions, insects and sessions that are embedded in the same direction are more likely to be connected, and the ones in the opposite direction are less likely to be connected. On the contrary, insects and sessions that are embedded in the same direction for the third and fourth dimensions are less likely to be connected, while the ones in the opposite direction are more likely to be connected. The choice of $D_{+}=D_{-}=2$ is justified by looking at Figure 15. The session-pollinator embedding can be seen in Figure 16 and Figure 17.
C.1 Spipoll, exploration with higher dimensional latent spaces
Figure 15: Estimated mean on 10 trials for the AUC(B^)AUC($\widehat${B}) (left) and $AUC(\widehat{B^{\prime}})$ (right) for link prediction in the Spipoll data set using BGVAE (top) and the fair-BGVAE (bottom) for various values of $D_{+}$ and $D_{-}$.
In Section 5 we show in detail the results for the case where the latent space has 4 dimensions with $D_{+}=D_{-}=2$. We justify this choice by looking at the estimated mean of the AUC(B^)AUC($\widehat${B}) for different numbers of dimensions for the latent space. Looking at Figure 15, we can see that the AUC(B^)AUC($\widehat${B}) doesn’t significantly change for higher values of $D_{+}$ and $D_{-}$.
C.2 Latent space representation